A fly flies with velocity v(t)=|10-t | m/s for 20 sec. What will be the travelled distance and average speed of the fly in 20 sec 

Grade:11

3 Answers

Saurabh Koranglekar
4 years ago
Dear student

Distance traveled per second = { 9,8,7,6,5,43,2,1,0,1,2,3,4,5,6,7,8,9,10}
Sum = 55+45= 100 m
Avg speed = 100/20 = 5 m/s

Regards
Saurav
4 years ago
i dont think you get that much time to solve as done by Saurabh Korangle
see the velocity time graph of above equation is a straight line so you should go by graph method
 graph is straight line (in particular :- two straight lines ) that is v=10-t for t10
finding area of two triangles (base 0 to 10 and base 10 to 20) area = 50+50=100m=total distance=total displacement as velocity is always positive in this case
average spped = total distance/total time = 5m/s 
it may appear longer but its faster
 
aprove if it helps you as it motivates me to answer more   :)
 
Saurav
4 years ago
typo in 3rd line v=10-t till t=10s and v=t-10 for t>10s
 
 
sorry for the typo
forget not to aprrove and feel free to ask any further questions